OpenStudy (anonymous):

find the slope of a line passing through the given points (-4 2) and (1 -3)

jimthompson5910 (jim_thompson5910):

Slope: m = (y2-y1)/(x2-x1) m = (-3-2)/(1-(-4)) m = (-3-2)/(1+4) I'll let you finish up

jimthompson5910 (jim_thompson5910):

Keep in mind that (x1, y1) = (-4, 2) and (x2, y2) = (1, -3)